### v2

# A NICER view of the â edge-on pulsar PSRÂ J06143329

Abstract
Four neutron star radius measurements have already been obtained by modeling the X-ray pulses of rotation-powered millisecond pulsars observed by the Neutron Star Interior Composition ExploreR (NICER). We report here the radius measurement of PSRÂ J06143329 employing the same method with NICER and XMM-Newton data using Bayesian inference. For all different models tested, including one with unrestricted inclination prior, we retrieve very similar non-antipodal hot region geometries and radii. For the preferred model, we infer an equatorial radius of km for a mass of (median values with equal-tailed credible interval), the latter being essentially constrained from radio timing priors obtained by MeerKAT. A more complex model, fitting the data equally well, resulted in a consistent inferred radius. We find that, for all different models, the pulse emission originates from two hot regions, one at the pole and the other at the equator. The resulting radius constraint is consistent with previous X-ray and gravitational wave measurements of neutron stars in the same mass range. Equation of state inferences, including previous NICER and gravitational wave results, slightly soften the equation of state with PSRÂ J06143329 included and shift the allowed mass-radius region toward lower radii by m, which is compatible with previous analyses to within less than one standard deviation.
1 Introduction
Neutron stars (NSs) are known to harbor ultra-dense matter in their cores (up to ), whose microscopic state can be constrained by the macroscopic properties of NSs. The Neutron Star Interior Composition ExploreR (NICER, gendreau_neutron_2016) is a fast X-ray timing mission, whose primary goal is the measurement of the radii of fast rotating NSs, known as millisecond pulsars (MSPs), to yield constraints on their inside matter.
MSPs are produced through a recycling process, where an NS accretes matter from a companion star, inducing its spin-up to millisecond rotation periods via transfer of angular momentum (alpar_new_1982). Such accelerated MSPs have a magnetic field strong enough (G) to cause the acceleration of charged particles in the magnetosphere, which results in complex and pulsed multi-wavelength emission as the NS rotates. In particular, the magnetic field induces an electron-positron particle shower near the magnetic poles of the NS surface. This phenomenon deposits energy at the magnetic polar caps of the NS, which locally heats the surface to K, consequently emitting thermal radiation in the soft X-ray (ruderman_theory_1975; arons_pair_1981; harding_pulsar_2001; harding_pulsar_2002) observed as pulsed emission as the NS rotates.
Measurements of MSPs radii are carried out by estimating their compactness through the imprint of their strong gravitational field on the pulse profile from the aforementioned surface emission. Contrary to other kinds of NS emission (e.g., radio, non-thermal hard X-rays or -rays), such surface emission is strongly affected by spacetime curvature due to the vicinity of the emission region to the center of the gravitational potential (km). Other relativistic effects, such as Doppler boosting and angle aberration, also carry information on the radius. In practice, pulse profile modeling (PPM) is used to model these effects and to infer the source parameters, including its radius, and their uncertainties using statistical methods (pavlov_mass--radius_1997; zavlin_soft_1997; bogdanov_nearest_2013; watts_constraining_2019; bogdanov_constraining_2021). This method has now been extensively tested and used on four different MSPs observed by NICER, with different datasets (riley_nicer_2019; riley_nicer_2021; miller_psr_2019; miller_radius_2021; salmi_radius_2022; salmi_atmospheric_2023; choudhury_nicer_2024; dittmann_more_2024; salmi_radius_2024; salmi_J1231_2024; vinciguerra_updated_2024; hoogkamer_cross-comparison_2025). Such methods enable precise modeling of the surface emission at the cost of high computation requirements.
Mass and radius measurements of NSs lead to constraints on the equation of state (EOS) of dense nuclear matter by informing about its stiffness (Watts:2016uzu; Chatziioannou:2024tjq). Around nuclear densities, the EOS is informed by chiral effective field theory (EFT, see, e.g., Hebeler2013; Tews:2012fj; drischler_chiral_2019; Keller2023; Tews:2024owl). The electromagnetic wave measurements obtained with PPM, along with gravitational wave (GW) measurements (GW170817, abbott_gw170817_2018; GW190425, abbott_gw190425_2020), currently favor EOSs on the stiffer side of a EFT informed prior distribution (e.g., Raaijmakers21; rutherford_constraining_2024; Koehn:2024set). While nucleonic EOS are the simplest explanation, more exotic EOS families also exist (see lattimer_equation_2016; oertel_equations_2017; burgio_neutron_2021 and references therein for examples). However, none of these EOS models are significantly favored by observations, which highlights the need for more numerous and tighter observational constraints.
The MSP analyzed in this study is PSRÂ J06143329, an intermediate mass MSP rotating at Hz discovered from its radio pulsations (ransom_three_2011). It has also been found to exhibit soft X-ray pulsations, likely of thermal origin, with no sign of non-thermal pulsations as all the pulsed signal was below keV (guillot_nicer_2019). A white dwarf (WD) companion with a hydrogen atmosphere has been discovered (bassa_cool_2016), allowing for distance estimates using the photometric temperature of the companion. This distance was later confirmed with radio parallax (shamohammadi_meerkat_2024). Tight constraints on the mass and orbital inclination also exist from strong Shapiro delay caused by the edge-on view of the PSRÂ J06143329 system (reardon_gravitational-wave_2023; miles_meerkat_2025). Such prior knowledge greatly decreases possible degeneracies with PPM and provide important insights to yield a tight and robust radius measurement, which proves more complicated without prior mass and inclination information (vinciguerra_updated_2024; salmi_J1231_2024). PSRÂ J06143329 has been observed by both NICER and the X-ray Multi-Mirror Mission (XMM-Newton, XMM hereafter), with a deep NICER accumulated exposure of Ms. It is hence a good candidate for a new mass-radius measurement using PPM.
In Section 2 we describe the X-ray data used for this analysis and its extraction. Section 3 outlines the modeling procedure adopted throughout this analysis. In Section 4 we present the results from the inference runs. In Section 5 we assess the validity of our exploratory runs. In Section 6 we discuss the results and their implication on the EOS of dense matter. We conclude in Section 7 and summarize our findings.
2 X-Ray Event data
This section summarizes the reduction of the X-ray data used for inference. All the data products used for this analysis can be found in the Zenodo repository doi: 10.5281/zenodo.15603405 (zenodo).
2.1 NICER
This analysis of PSRÂ J06143329 makes use of 1.6âMs of raw NICER data acquired in the 2018 March 3rd â 2023 May 15 period (OBSIDs to ), comprised of all the observations of this source antecedent to the NICER light leak111https://heasarc.gsfc.nasa.gov/docs/nicer/analysis_threads/light-leak-overview/. The few available observations dating after the light leak occurrence (19âks) are characterized by a high noise level caused by optical loading and were not included due to the difficulty of properly screening them alongside their short additional exposure.
The screening procedure, to remove time intervals with high background, generally follows the one adopted in Bogdanov2019a and by the other NICER PPM analyses listed above. Calibration was handled by the Heasoft software (2014ascl.soft08004N), version 6.33.2, which includes NICERDASv12 along with the calibration file xti20240206. Initial screening was produced with the nicerl2 task. Unlike previous analyses, no âhotâ focal plane modules (FPMs) were consistently screened out since, as part of the niautoscreen task, they are now individually removed when exhibiting count rates above the median of active FPMs. The final ancillary response file (ARF) is automatically scaled down accordingly, based on the average number of FPMs activated during the whole exposure. This allows to extract more events while limiting background.
Good time intervals (GTIs) were further selected with the psrpipe script from NICERsoft222https://github.com/paulray/NICERsoft, by removing time intervals which included any of the following:
âą
High optical photon flux causing considerable FPM reset: median undershoot rate cts/s.
âą
High particle flux causing excess high-energy detections: overshoot rate cts/s/FPM.
âą
Strong solar activity, when the geomagnetic storm index ( index) is high: .
âą
Spacecraft in a region of low shielding to particle flux, i.e., with low cut-off rigidity:
COR_SAX GeV/. -
âą
Strong remnant background after having applied previous criteria: count rate cts/s for s bins in the keV band. As we do not expect any source count in this energy range, this is a good proxy for background count rate.
These values are chosen to be similar to previous NICER MSP analyses. This procedure resulted in a final filtered dataset with Ms of effective exposure time.
Extracted events were phase-folded with the PINT task photonphase (luo_pint_2021), using a joint Nançay Radio Telescope (NRT) and Fermi timing solution spanning the entire NICER observation window, to produce a phase-resolved spectrum (Figure 1). The resulting bolometric pulse profile has a strong main peak and a weaker secondary peak, which is consistent with previous findings (guillot_nicer_2019). Finally, the associated ARF and response matrix file (RMF) were extracted using nicerl3-spect task.
The analyses were performed on pulse invariant (PI) energy channels , corresponding to an energy band of keV. While the lower energy bound is motivated by the NICER effective area sudden drop as well as residual optical photon flux background below keV, the upper bound is motivated by the low source flux and absence of pulsed emission over keV, making the total flux background dominated above this energy. A Kolmogorov-Smirnov test comparing the phase distribution of photons with energy over keV to a uniform distribution did not find any significant deviation (). All inferences pointed toward a contribution of the source to the total NICER flux lower than at keV (see Section 4), validating the choice of this upper limit.
The NICER phase-averaged spectrum shows two prominent features near keV and keV respectively (see Section 4.4). Both are phase-independent peaks, the latter appearing mostly in observations after mid-2022. We identify it as a fluorescence line of oxygen VII at eV in the Earth atmosphere, which is excited by the stronger solar activity at later dates. This line also already appeared prominently in NICER spectra of some analyses (salmi_radius_2024; salmi_J1231_2024) and subtly for other analyses that did not include any observations after the solar activity started rising (mid-2022), and for which the source flux was higher (choudhury_nicer_2024; vinciguerra_updated_2024).
MSP surface emission models cannot explain such time-independent narrow spectral features. Hence, such an excess is accounted for as a phase-independent background, which is fitted alongside the modeled pulse profile (more details in Section 3.5). We ensured that these spectral features were well reproduced by the fitted phase-independent background by inspecting the inferred background for all runs. Such peaks being also inferred as background in previous analyses further supports that PPM accounts for them in the phase-independent background, even in cases where the O VII line was less prominent (vinciguerra_updated_2024).
Although NICER background models such as Space Weather (SW) or 3C50 (remillard_empirical_2022) exist, none has been used in this study and the NICER background is inferred directly from the data. This approach is facilitated by the available XMM observations, which inform on the phase-averaged source spectrum and flux, which in turn inform the NICER expected source spectrum, and hence its background spectrum. salmi_radius_2022 has shown that including background constraints from XMM for PSRÂ J07406620 gives comparable results to using SW or 3C50 without relying on any of the assumptions that either of them makes. Moreover, 3C50 relies on observations from before mid-2020, when the OÂ VII line was not as prominent, which can lead to underestimation of the contribution of this background component to the flux. As one XMM observation was available, we decided to use it instead of one of the background models.
2.2 XMM-Newton
The XMM data for this source consist of a single observation taken on 2010 October 4 (OBSID ) with the imaging mode. All EPIC instruments were used (PN, MOS1, MOS2), each providing phase-averaged source and background spectra. These were extracted using the XMM Science Analysis Software SAS version 21.0.0. The data were screened with the recommended filters (FLAG and PATTERN for PN and PATTERN for MOS1/2). No flaring particle background occurred during the observation and the entire exposure was kept. This resulted in a final 11.36Â ks effective exposure for PN and 14.42Â ks for MOS1/2.
The source spectra were extracted from a 35âł circular region, while the backgrounds were extracted from a large 35-550âł annulus region centered on the source, and limited to the same charge-coupled device. Other X-ray sources were excluded and low vignetting regions avoided. The exact regions differ between EPIC instruments. The backscal parameter and respective exposure times written in the processed observation files were used to scale the background spectra to the sources spectra. The ARF and RMF were extracted using the arfgen and rmfgen tasks.
For our analyses, we used the XMM counts within the same keV energy band as for NICER (Figure 2). Moreover, the observed source counts fall to background level at around 2.0âkeV, making the flux above this energy undistinguishable from background.
3 Modeling
The modeling procedure as a whole is similar to that of riley_nicer_2021; salmi_radius_2022; choudhury_nicer_2024; salmi_radius_2024; vinciguerra_updated_2024. This section summarizes the PPM, the parameter inference method and adopted settings. All relevant files for modeling and inference are provided in the Zenodo repository (zenodo).
3.1 X-PSI settings
We model the energy dependent pulse profile using the X-ray Pulse Simulation and Inference package (X-PSI, xpsi) v2.2.7. For post-processing of the inference runs only, we used X-PSI v3.1.0.
The resolution settings of X-PSI (see Appendix A for explanations on the different settings) have below 1% effects on the computed waveform for simple geometries and can reach a few percent under extreme cases with thin and elongated spot structures (choudhury_exploring_2024). Due to the limited count statistics of our dataset, we deemed such complex geometries to be difficult to distinguish from simpler ones. Hence, these were not investigated during our exploration runs which were performed using the low resolution (LR) X-PSI settings from choudhury_exploring_2024. This allowed a lower computational cost and the use of a better sampling resolution instead. However, our headline result uses higher X-PSI resolution to compute accurate waveforms for those extreme cases and mitigate the effects of a limited X-PSI resolution on the sampling procedure (see Appendix A). These effects are further investigated in Section 5.
3.2 Sampling
The posterior distributions of model parameters were obtained with nested sampling (skilling2004nested). Among other Bayesian inference algorithms, nested sampling has the advantage of providing thorough sampling over a large prior parameter space and easy model comparison with limited number of hyper-parameters. In practice, sampling was carried out using PyMultiNest (PyMultiNest), a Python wrapper of the nested sampling implementation MultiNest (MultiNest_2009). Although UltraNest (buchner_ultranest_2021) is in theory more robust, it is also much slower and performed equally well as MultiNest for PPM on PSRÂ J07406620 (hoogkamer_cross-comparison_2025). We hence opted to use MultiNest for its lower computational cost.
All exploratory runs referenced here used the same set of hyper-parameters, chosen to ensure good sampling of the parameter space, while limiting computation time and allowing for easy and fair comparison between runs. MultiNest has three hyper-parameters, which were chosen to be live points (LP), a sampling efficiency (SE) of and an evidence tolerance (ET) of for the exploratory runs (see MultiNest_2009 for a description of these hyper-parameters). We ran all inferences with multi-modal exploration enabled. This allocates distinct sampling regions to the different modes identified, allows for individual post processing of the latter, and may speed-up the computations (MultiNest_2009). Such hyper-parameters grant a more complete sampling (for exploratory runs) than in previous studies. Our headline result was run with LP, SE and ET, which allows for a higher sampling accuracy. The exploration of the parameter space completeness is further discussed in Section 5.
3.3 Priors
Unless stated otherwise, we settled on the following common prior configuration for all exploratory runs. The prior equatorial radius samples are drawn from a uniform distribution km, with the surface gravity and compactness conditions rejecting samples with high compactness (see Appendix B). This prior is different than those in previous analyses, which drew samples down to km.
bassa_cool_2016 discovered a WD companion to PSRÂ J06143329. Its observed temperature from optical photometry resulted in an estimated distance in the range âpc, consistent with radio timing parallax measurements (âpc, shamohammadi_meerkat_2024). For our inference, we used a distance prior following a normal distribution , corresponding to the optical photometry constraint as a range, truncated at .
Tight constraints on mass and inclination were obtained by MeerKAT from radio timing thanks to the detection of a strong Shapiro delay (miles_meerkat_2025). These resulted in a mass and orbital inclination of, respectively,333These values are derived from the radio solution provided in the -years MeerKAT PTA data release. and , used in our inference as Gaussian priors, with the sine of the inclination truncated at one. These are consistent with the measurements from reardon_gravitational-wave_2023. As in previous X-PSI analyses, we considered the spin axis of PSR J06143329 to be aligned with its orbital momentum, which is expected to be the case at the end of the recycling process (Yang_inclination_2023). The unconstrained inclination hypothesis was also tested and resulted in similar posterior distributions (see Section 4.2).
The interstellar absorption was modeled with the Tbabs444https://pulsar.sternwarte.uni-erlangen.de/wilms/research/tbabs/ model, depending only on the neutral hydrogen column density parameter . The value from hi4pi_collaboration_hi4pi_2016 in the direction of PSRÂ J06143329 is . This value is possibly overestimated as it is integrated along the whole line of sight. Using the relation between radio dispersion measure (DM) and from he_correlation_2013 with yields , but this relation is empirical and has a large scatter associated with it. Since there are uncertainties on those measurements, we adopted a wide uniform prior of .
3.4 Atmosphere model
Knowledge of the atmosphere of the NS is also crucial for PPM, as the inferred radius and geometry can depend on the atmospheric composition (salmi_atmospheric_2023). The hydrogen atmosphere of the DA spectral type WD companion, from which the NS likely accreted material to spin-up, heavily favors the hydrogen atmosphere hypothesis. Because of gravitational settling, the lighter elements rise to the top of the atmosphere, making the accreted hydrogen the only visible component of the NS atmosphere. Using a model with a different composition implies the assumption of a total depletion of hydrogen in the NS atmosphere. This could in principle happen due to diffuse hydrogen burning, but the timescale for this process to complete is uncertain (chang_diffusive_2003; chang_evolution_2004). We chose to use a fully ionized NSX hydrogen atmosphere (ho_atmospheres_2001) for our exploratory runs, as done in all our previous analyses. A single inference run with a partially ionized NSX atmosphere table (used previously in miller_radius_2021) was also carried out for our most favored model (see Section 4.2).
3.5 Background
A phase-independent background is inferred by fitting a background parameter for every instrumental energy channel. Priors of these parameters can be informed by background measurements, models, or left free to vary from up to the total number of observed counts when a more precise prior value is not available. However, to save on the computational expense from sampling hundreds of supplementary parameters, these parameters are marginalized rather than being sampled directly. The procedure is as follows: for a set of source parameters, X-PSI computes the expected source phase-resolved spectrum and a supplementary contribution to the otherwise Poisson likelihood is evaluated. This supplementary contribution is based on the marginalization of background parameters over the background prior, which is computed in practice within a region of the parameter space around the best fitting background parameters (for a precise derivation of this contribution, see riley_neutron_2019, Appendix B.2). These best fitting background parameters associated with the source parameters are the background parameters used for plotting in Section 4.4. This method allows for much faster computations.
Previous works made use of 3C50 background estimates as priors for the NICER background (salmi_radius_2022; choudhury_nicer_2024), but we chose here to leave it uninformed and use phase-averaged XMM data instead. The information from the phase-averaged source spectrum derived by XMM allows X-PSI to infer constraints on the source spectrum in the NICER data, hence resulting in NICER background constraints without relying on 3C50 assumptions (see Section 2.1).
The same background marginalization procedure is applied to the phase-averaged XMM data. However, contrary to NICER, the background prior for XMM is tighter as it is informed by the data. As in previous analyses (riley_nicer_2021; salmi_radius_2022; choudhury_nicer_2024; salmi_radius_2024; vinciguerra_updated_2024), the backgrounds of the different EPIC instruments were left to vary uniformly in a range around the measured EPIC background spectra, where is the uncertainty on the number of counts generated by a Poisson process (Figure 2). We adopted throughout all our analyses, whereas previous analysis of PSR J07406620 with XMM data used a value of to account for possible biases in the background, which was not extracted from the same exposures but rather from blank-sky exposures (salmi_radius_2022; salmi_radius_2024). The XMM background spectra for this analysis were extracted from the same exposures as the source spectra, hence this constraint was tightened. This did not prove problematic as, for the EPIC instruments, the contribution of the background to the flux (âcts/bin) is negligible compared to Poisson uncertainty (Figure 2).
3.6 Instrument responses
We ran joint NICER and XMM fits for all our inference analyses, hence the cross-calibration uncertainties between the instruments had to be taken into account. Each instrument ARF was multiplied by an energy-independent fitted scaling factor parameter. These parameters are , , , , and were all fit using the same uncorrelated Gaussian prior . This choice relaxes assumptions adopted in previous analyses, which used a single factor for all XMM instruments.
An exploratory run with uniform priors instead, , resulted in broader scaling factor posteriors still centered around . Interestingly, it showed a strong positive correlation across XMM instruments and no visible correlation with NICER. However, maximum likelihood cross-calibration scaling factors pointed toward the edge of the prior distribution, which is highly unlikely based on current knowledge of the uncertainty on cross-calibration between instruments (, marshall_concordance_2021), hence resulting in biased best fit values. Hence, we opted to keep the Gaussian priors of the scaling factors for all inference runs.
4 Inference
In this section, we present inference results for the different hot region geometry models that have been tested, and their respective concordance to the data. The taxonomy of different models is the same as in riley_nicer_2019. Results and posterior distributions for all the models can be found in the Zenodo repository (zenodo).
4.1 Exploratory ST-U
Model definition â The observed double peaked bolometric pulse profile hints at the presence of at least two hot regions. A simple ST-Uâ model (Single Temperature - Unshared) was tested. It is composed of two non-overlapping circular hot regions with independent positions, sizes, and temperatures.
Geometry and modes â The inference analysis results in two modes with similar configurations of the hot regions, where one spot is located near the equator while the other is on either of the rotational poles. This arises from the edge-on viewing geometry of PSRÂ J06143329, which results in a North/South degeneracy on the location of the polar hotspot. Such an effect was already observed during the analysis of PSRÂ J07406620, which is also seen edge-on (riley_nicer_2021; salmi_radius_2022; salmi_radius_2024). Both modes have highly similar () marginal 1D posteriors on non geometric parameters, are equally likely, and neither is favored by the data.
NICER Background â The phase-independent inferred background spectrum has features at the energies of the peaks identified in the data (see Section 2.1). As expected, X-PSI picked up these contributions as part of the phase-independent background. The inferred spectrum of the NS is made of two thermal components and has no peak around this energy. The background ranges from 75% of the signal at low energies up to of the signal at keV, hence the chosen upper energy limit of the fitted data is well suited.
Inferred parameters â The main peak in the NICER pulse profile (see Section 2.1) at phase (Figure 1) is caused by the equatorial spot, going in and out of view during rotation. The weak peak is caused by the polar hotspot at , which never fully goes out of view. The phases and temperatures of both hotspots are well constrained for both modes (respectively, and ). Spot sizes are only constrained to about . The colatitude of the equatorial hotspot is not tightly constrained (). Finally, the radius is rather small and tightly constrained, with km. The full ST-Uâ posterior is available in the Zenodo repository (zenodo).
Model performance metrics â Values of log-evidence and maximum log-likelihood are given in Table 1. The 2D residuals of all the tested models can be found in the Zenodo repository (zenodo). ST-Uâ can reproduce the data as the residuals do not feature any significant deviation from random Poisson noise. However, we tested more complex hotspot geometries to check whether it would bring any improvements.
| Model | ||
| ST-U | ||
| ST+PDT | ||
| PDT-U | ||
| ST+PDTâ - | ||
| ST+PDTâ - H | ||
| ST+PDTâ - | ||
| ST+PDTâ - | ||
| ST+PDTâ - Restricted prior | ||
| ST+PDTâ - Headline |
4.2 Exploratory ST+PDT
Model definition â We now explore the ST+PDTâ (Single Temperature - Protruding Double Temperature) model. Compared to ST-U, this model keeps one circular single temperature region (the ST region), while the second hotspot is replaced by two overlapping circular spots, each with its own non-null temperature (the PDT region). The spot being overlapped is referred to as ceding and the overlapping one as superseding. A PDT spot has four more parameters than a ST spot. They parametrize the ceding spot size, relative position to the superseding spot and temperature while the superseding spot is parametrized like an ST hot region.
Geometry and modes â The inference resulted in four modes, all with a geometry similar to that of ST-U: one hot region at the equator and another at the pole, with a North/South degeneracy. Although the respective colatitude of the ST and PDT regions is not constrained a priori, the new PDT region ends up being always located on either pole, while the ST region is always at the equator.
The ST region parameters have similar posterior distributions as the equatorial ST region in the ST-Uâ model, with a well constrained phase (), size (), and temperature (), but a poor colatitude localization ().
The PDT region has two possible configurations that explain the data equally well, each with a North/South degeneracy. While both PDT configurations are made of a large and rather cold region (with angular radius and temperature K) along with another small and hot region ( and K), these regions can be either ceding or superseding components. In both configurations, the small and hot regions arrange themselves on the surface to approximately reproduce a similar surface and position of the hotspot as the polar hot region from the ST-Uâ model (see Section 4.4 where a visual representation of similar configurations is provided).
Compared to the posterior predictive ST-Uâ bolometric pulse, the sharpness of the main peak is slightly increased to better reproduce the data. However, the new PDT is not mainly responsible for this pulse and has settled into a configuration where it partly contributes to the main peak while being fully responsible for the weak peak.
NICER Background â The inferred background is similar to that of ST-Uâ as it reaches the registered NICER signal at keV and still features the peaks at âeV and eV to account for the phase-independent peaks.
Inferred parameters â The inferred radius of km is higher than for ST-Uâ by 1.2âkm, but is still compatible at (Figure 3). However, the radius values also depend on the mode, where the hot superseding configuration corresponds to a radius lower than the hot ceding configuration by 0.5âkm. Comparing mass-radius posteriors, it seems that the ST+PDTâ distribution corresponds to the high radius tail of ST-U. However, this tail was not classified as a distinct ST-Uâ mode by MultiNest. The full ST+PDTâ posterior is available in the Zenodo repository (zenodo).
Model performance metrics â The ST+PDTâ residuals do not feature any deviation from random Poisson noise and is able to explain well the data. The log-evidence of this model is decisively better than the value of the ST-Uâ model using the Kass and Raftery criterion (kass_bayes_1995, here ). Moreover, the max-likelihood also shows strong improvement. All values are reported in Table 1. This makes ST+PDTâ a more suitable model than ST-Uâ to explain the data.
Supplementary runs â We relaxed the orbital angular momentum and spin alignment hypothesis (see Section 3.3) for one ST+PDTâ run by leaving the inclination angle unrestricted. It resulted in three modes, each replicating distinctive geometric spot configuration among the four modes that were seen in the exploratory ST+PDT, but this time with slightly lower inclinations. All modes have inclination and radius constraints consistent with the reference exploratory ST+PDTâ model, with respective inferred values of , and . This indicates that the assumption of alignment of the spin axis with the orbital momentum is not problematic in this case. This run was more computationally expensive while resulting in similar, although not as tight, radius and geometry constraints. Hence, this configuration was only tested for this model.
To assess the effects of a different atmosphere model, an alternative ST+PDT exploratory run was carried out using the partially ionized NSX hydrogen model (see Section 3.3 for the choice of this atmospheric model). We retrieve the same four geometric modes as for the fully ionized atmosphere case with slightly higher, but compatible, equatorial radius constraints (with difference in the median values km). The log-evidence with the partially ionized atmosphere model is better, but not significantly ( kass_bayes_1995). Hence this atmosphere model gives results consistent with those of the fully ionized model, which we use for the rest of the inferences.
4.3 Exploratory PDT-U
Model definition â A PDT-Uâ model (Protruding Double Temperature - Unshared) is composed of two PDT hotspots. Although the ST+PDT results were convincing, it was tested to investigate the possible metric gains that a second PDT hotspot would bring. Note that this model is a much more computationally intensive one than ST-Uâ and ST+PDT.
Geometry and modes â The inference results in two configurations, each with two North/South degenerate modes, still with one region at the pole and the other at the equator. As for ST+PDT, the PDT-Uâ model has a configuration where the ceding region is hotter and smaller than the associated superseding one. However, this region is now located at the equator while it was the polar region for ST+PDT. Moreover, the geometric parameters posterior distributions are still similar to those of both ST-Uâ and ST+PDT. The polar hot region, in particular, is fully consistent with the ST+PDTâ posterior in the hot superseding configuration (see Section 4.2 for more details on this configuration).
NICER Background â The inferred background of this model is, again, similar to that of ST-U. It still shows an excess around the peaks in the spectrum and reaches the registered NICER signal at âkeV.
Inferred parameters â The inferred PDT-Uâ radius km is compatible with the ST+PDTâ one, with a difference in median values of (, Figure 3), which makes them consistent with one another.
Interestingly, for PDT-U, all individual modes yield a higher and than the combined distribution. This was not the case for ST-Uâ or ST+PDT. This happens when the distributions of the different modes intersect only for lower values of and in some parts of the parameter space. Such a phenomenon was already observed when combining inference results from different bursts in kini_constraining_2024.
The posterior distribution of the angular radii of both superseding hotspots spans rather low values around , with the best-fit parameters reaching even lower values with sizes of across both superseding hotspots in both modes. The full PDT-Uâ posterior is available in the Zenodo repository (zenodo).
Model performance metrics â Although PDT-Uâ has a better log-evidence and maximum likelihood than ST+PDTâ (, Table 1), indicative that PDT-Uâ performs better, such a small difference is not decisive according to the Kass and Raftery criteria (kass_bayes_1995). vinciguerra_x-psi_2023 previously found that a log-evidence difference of a few units could not tell models apart: in that work, the log-evidence of a more complex model was units greater than for a simpler model, although the data were generated by the simpler model. It is a possibility that the evidence of PDT-Uâ increases, although not significantly, by accumulating evidence in the region of the parameter space where it reproduces ST+PDT.
Finally, PDT-Uâ is much more computationally expensive to run, while resulting in similar geometric and radius constraints, which are our main parameters of interest. Therefore, for our high resolution run we opt to continue with the simpler ST+PDTâ model, which already reproduces well the data. Further explanation on this choice is given in Section 5.
4.4 Headline ST+PDT
Model definition â The Headline run was carried out with the ST+PDTâ model, with improved X-PSI resolution settings and sampling hyper-parameters (see values in Appendix A and B) and still with the multi-modal variant of MultiNest. Because the omnipresence of the North/South degeneracy across previous runs, which gave similar radius constraints, slowed down the computation times and made thorough sampling harder, geometric priors were constrained to remove this degeneracy. This was effectively done by forcing the polar PDT region to lie on the North hemisphere of the NS. A test run with the exploratory ST+PDTâ showed that this procedure speeds up computation by a factor of 2 while resulting in similar geometric and radius constraints, with a difference of radius of âm, the same order of the âm Monte Carlo standard error, an uncertainty inherent to sampling procedures. Moreover, restricting the prior volume allows for more thorough sampling, making it harder to miss high likelihood regions.
The choice of the North hemisphere, instead of the South hemisphere, for the location of the PDT hot region allows for more conservative uncertainties on the radius. This is because the two North PDT configurations of the exploratory ST+PDTâ have individual median radii that are, respectively, the lower and higher values among all the modes found in this exploratory run, with uncertainties greater or equal to that of their southern counterpart.
Geometry and modes â This analysis uncovered four modes, two of which were already previously found in the exploratory run. The two new modes both having significantly lower log-evidence () are hence disfavored and have only a barely visible impact on the total posterior distribution. These modes also appeared in the run testing the setup to break the degeneracy and are not discussed here, but their presence itself reveals that MultiNest finds even low evidence modes. The two modes existing in the headline run have nearly the same distributions as previously found in the exploratory ST+PDTâ run. The best fit geometry of each mode is shown on Figure 4. Analogous to the exploratory ST+PDT, the PDT spot is solely responsible for the weak peak and partially contributes to the main peak (Figure 5).
NICER Background â The background is similar to the one of exploratory runs. However, the deviation of the maximum likelihood background from the inferred distribution median is 3 at energies below âkeV, and is decreasing with increasing energy (right panel of Figure 5). However, note that we do not fit the individual ST, PDT, or background components individually but their sum and the best fit geometry are consistent with the total spectrum expected distribution, which also has smaller uncertainties than individual ST, PDT or background contributions (Figure 5).
The uncertainties in the contributions of the different components arise from uncertainties in the NICER source spectrum, which is mostly dependent on the XMM spectrum and cross-calibration factors (see Section 3.5). As the latter are poorly constrained and the count statistics of the XMM observation is low, the accuracy of the background constraint is limited.
In practice, the best fit parameter vector of this headline ST+PDTâ run has ARF scaling factors (see Section 3.6) lying on the edge of the prior distribution. This is compensated by the background marginalization, resulting in a background spectrum on the lower edge of the distribution along with a high likelihood value.
Hence, although it is consistent with the posterior geometric distribution, the best fit geometry should be considered with care. Whenever possible, the full posterior distributions of the geometric parameters should be used instead. The maximum a-posteriori parameter vector of the individual modes could also be considered. These can be retrieved in Appendix B (for the geometric parameters) or from the Zenodo repository (for the full parameter vector, zenodo).
Inferred parameters â For this headline run, we infer the radius of PSRÂ J06143329 to be âkm for a mass of . This is consistent with the exploratory ST+PDTâ run and both exploratory ST-Uâ and PDT-Uâ runs.
The inferred distribution of reaches the lower bound of the prior (see Appendix B for the posterior distribution of a subset of parameters, the full Headline ST+PDTâ posterior is available in the Zenodo repository from zenodo). This value is compatible with the hi4pi_collaboration_hi4pi_2016 value, the latter being also compatible with .
A diagnostic to test the convergence of MultiNest is to see whether increasing sampling accuracy affects the posterior distribution and evidence (higson2019diagnostic). Here, the posterior is similar to that of the LR run with the same restricted priors, although uncertainties are slightly larger. This indicates that sampling has likely converged.
Model performance metrics â As evidence is proportional to the prior, which was restricted for this Headline run in order to remove the North/South degeneracy, the new log-evidence can not be directly compared to the exploratory run value. Instead, it should be compared to the exploratory ST+PDTâ run with restricted priors value (see Table 1), whose log-evidence is consistent within with the Headline value.
The posterior predictive bolometric pulse (see left panel of Figure 5) fits well the data. The PDT spot has higher flux and more uncertainty associated to its inferred spectrum than the ST spot (see right panel of Figure 5). This is not due to the multi-modality and happens for both configurations, indicating that the PDT parameters are more difficult to constrain than the ST ones in our case.
5 Validation
In this section, we describe the different steps taken to assess the validity of the results from our exploratory and Headline runs.
5.1 X-PSI resolution
As we used LR settings for the exploratory runs and mixed resolution (XR) settings for the Headline run (see Appendix A for more information), we investigate the effects of the settings configuration on sampling.
The effects of the LR settings on the likelihood were investigated by drawing samples from the posterior of the models presented in Section 4 and computing a likelihood with high resolution (HR) settings for each sample, which was then compared against the likelihood with LR (exploratory) or XR (headline) settings used during sampling. The distributions are plotted in Figure 6. The distributions are all compatible with at the level. The spread, which measures the accuracy of the likelihood, increases with model complexity, reaching for PDT-U. This is because, under the same X-PSI resolution settings, we observe a higher likelihood accuracy for ST regions compared to PDT regions, as the former exhibit a simpler geometry. While the accuracy (i.e., a maximum value of ) needed for precise likelihood computation is difficult to determine, we deem these values low compared to the range of values taken by the likelihood during sampling, . Moreover, a log-evidence difference of would be needed to claim a significant superiority of a model over another, which is times more than our reported accuracy. Hence we could not find any systematic bias between LR and HR likelihoods for the posterior geometries, although we can conclude that PDT-Uâ likelihoods were less accurate than those of the other models.
Even if their likelihoods under HR settings are higher, configurations with more complex, thin, elongated structures with biased LR settings likelihoods could be missed while sampling. However, these distributions of LR runs support LR being accurate in the posterior space of geometries for all ST-U, ST+PDTâ and PDT-U, and their posteriors are consistent with a local maximum of the HR model. Nevertheless, we did not demonstrate that LR settings yield accurate likelihoods over the whole prior space of geometries and it is hence not possible from such considerations to conclude that a complex structure was not missed. It is possible that, somewhere in the prior space, there is a better fitting geometric configuration for which LR likelihoods are not accurate and biased toward lower values. Such a region could be unknowingly missed during sampling and would appear when using high X-PSI resolution. The Headline ST+PDTâ model was run with a higher X-PSI resolution and sampling accuracy to account for this possibility and converged toward the same solution as the exploratory run (see Section 4.4). This indicates that it could also be the case for PDT-U, which can be confirmed with a higher resolution PDT-Uâ run.
For the Headline ST+PDTâ model, we find a Gaussian-like distribution of with and (Figure 6). As these values are one order of magnitude lower than for LR runs, we deem this resolution increase sufficient. Moreover, it would be too computationally expensive, given current computational resources, to further increase the resolution (see Appendix  A).
5.2 Noise realization
The Poisson noise in the NICER data can have an effect on the inferred parameters (vinciguerra_x-psi_2023). This phenomenon has also been investigated here by testing different binning of the phases for the PSR J06143329 data. Even though the phase measurement of all collected photons is very precise (s), the data used for likelihood computation (Figure 1) is binned in rotation cycles (about s wide) bins to save on computational time. By uniformly shifting the phase of individual NICER photons by a small amount ( bin) before binning, most photons will remain in the same bins, and some will end up in the neighboring bin, therefore resulting in slightly different bolometric pulse profiles. This way, we mimic different noise realizations of the binned pulse profile, which, however, all share the same spectrum and XMM data. Using shifts of and cycles, that are smaller than the bin width, we ran LR runs with the preferred ST+PDTâ model.
All runs resulted in similar posterior distributions for all the parameters, except for the phase of the hot regions (Figure 7). As expected, the observed shift of the inferred hot region phases and matches exactly the shift applied to the data (although this is more difficult to visualize for the secondary hot region, as has two modes with large phase separation). However, one mode is missing for the cycles shift, which indicates that the parameter space is not fully sampled. On the other hand, no other mode was found during these analyses, which indicates that sampling likely did not miss any significant supplementary mode because of noise realization.
The log-evidences and maximum likelihoods have strong dependence on the binning of the data. This dependence (given in Table 1) is greater than the difference in metrics between all the ST-U, ST+PDT, and PDT-Uâ models.
This check allows to ensure that the solutions inferred by the previous exploratory runs are not artifacts from noise realization being over-fitted, but rather are consistent across different noise realizations.
5.3 Model comparison
As it was not trivial to differentiate ST+PDTâ and PDT-Uâ from log-evidence values alone (see Section 4.3), we turned to the arviz package (vehtari_practical_2017), which provides model comparison techniques taking into account the complexity of the models and comparing their predictive power. This requires a point-wise likelihood, i.e., a likelihood value for each individual bin. This product is, however, not available in our case, since our inferences use channel-wise background marginalized likelihoods, which is not applicable bin by bin (see Section 3.5). Instead, we use a simple Poisson likelihood and specify as supplementary input parameters the background parameter values.
Using samples, the variance of the expected log point-wise predictive density (ELPD) is large for all models, making them all compatible, but the variance of difference of ELPD is not as large (Figure 8). The significance of these deviations can be measured with two metrics (Leave-One-Out and Weighted Akaike Information Criteria), both of which favor PDT-Uâ over ST+PDT, but not significantly, while ST-Uâ is totally disfavored. This is in line with our previous conclusions using evidence. However, using the ELPD metric likely introduces some bias in our case, as the Poisson likelihood is different from the background marginalized likelihood used during sampling. As this model comparison could not clearly disentangle PDT-Uâ and the simpler ST+PDTâ model, which both result in similar radius constraints, and given the lower computational time of ST+PDT, we chose to use the latter for the headline model.
6 Discussion
6.1 Comparison to other MSPs with PPM constraints
The inferred radius of PSRÂ J06143329 is constrained at the level ( credible interval)555These numbers do not include systematic uncertainties, which are difficult to quantify. A rough estimate can be obtained by looking at the dispersion of radius values for different models. For instance, it is of the order of for PSRÂ J06143329 and for PSRÂ J00300451., which is not as tight as for PSRÂ J04374715 (about , choudhury_nicer_2024), but is of the order of PSRÂ J07406620 (about , salmi_radius_2024). The radius constraint of PSRÂ J12311411 (salmi_J1231_2024) is harder to compare, since that study used different radius priors, some of which were EOS informed, causing the tighter constraint. Some individual models for PSRÂ J00300451 achieved lower uncertainty thanks to higher photon statistics, resulting from a deeper exposure and a brighter source (riley_nicer_2019; vinciguerra_updated_2024).
While vinciguerra_updated_2024 incorporated background constraints in the PSRÂ J00300451 analysis, riley_nicer_2019 did not. The former resulted in different models fitting the data equally well, partially because of the lack of a mass prior. Our result is consistent with the ST+PDTâ solution for PSRÂ J00300451 from vinciguerra_updated_2024. However, our inferred radius is not directly comparable to that of the PDT-Uâ solution from PSRÂ J00300451, as this solution points toward a mass of 1.7â instead of 1.4â.
Interestingly, the measured radius is compatible at the level with the inferred radius of PSRÂ J04374715, a MSP of a similar mass (choudhury_nicer_2024) and with the indirectly inferred666GW events carry information on tidal deformability. Extracting radius information from this information is an EOS-dependent process. mass-radius contours for the NSs whose merger produced the GW170817 signal (abbott_gw170817_2018).
The PSRÂ J06143329 constraints presented in this work are deemed highly reliable as the inferred parameters are consistent across all different geometric models, for different X-PSI resolutions and different prior assumptions tested, such as the spin-orbital momentum alignment and atmospheric ionization state. These results all have a median radius in the âkm range. They all share a similar non-antipodal geometry, with one hot region on the pole, which is always partially visible, and a hot region near the equator, going in and out of view and causing the main peak in the profile.
This consistency across all runs in our analysis could be explained by the tight priors on mass and inclination. PSRÂ J00300451 and PSRÂ J12311411 are missing such constraints, making the inferred radius and mass ranges different from model to model. Although PSRÂ J07406620 has such constraints and times more exposure than PSRÂ J06143329, it is also twice as far and times fainter.
PSRÂ J04374715 is very bright with a tight prior on its mass, inclination and distance (reardon_neutron_2024) and also has little change in measured radius across models (choudhury_nicer_2024), but an AGN is in the field of view and a 3C50-informed background was used, both introducing possible bias. Moreover, PSRÂ J04374715 is seen at a inclination, which may allow for a completely different hotspot configuration when changing models. This does not seem to be the case for PSRÂ J06143329, which is seen edge-on. Overall, PSRÂ J06143329 likely owes its radius measurement stability to its tight prior from radio timing, along with the fact that it is seen edge-on.
6.2 Implications of PSRÂ J06143329 on the dense matter EOS
With the new PPM-derived mass-radius posteriors of PSRÂ J06143329 in hand, a subsequent EOS inference is necessary to understand the impacts of this new measurement on the inferred dense matter EOS and NS properties. To perform an EOS inference, we build on the work of rutherford_constraining_2024. Starting from the Baym-Pethick-Sutherland crust EOS (Baym71) for densities below , where is the nuclear saturation density, we use the EOS band from next-to-next-to-next-to-leading order (N3LO) chiral effective field theory (EFT) calculations up to (Keller2023). Beyond this density, we use two different high-density EOS parameterizations: one based on a piecewise-polytropic model (PP, Hebeler2013) and one based on the speed of sound in the NS (CS, Greif19). This results in a very large prior range for the EOS. For details on how the dense matter EOS models are constructed and their prior assumptions, we refer to rutherford_constraining_2024.
We obtain posterior EOS distributions through sampling with MultiNest. This entails first evaluating the priors based on allowed EOSs in the EFT range and for a broad range of parameters that span the PP or CS extensions. The likelihood incorporates the mass-radius constraints of MSPs obtained with PPM and mass-tidal deformability from GW events. This framework is implemented using the open-source dense matter EOS inference code NEOST777https://github.com/xpsi-group/neost (v2.1.0, Raaijmakers24). NEOST has been used for NICER EOS constraints and other dense matter inferences with prerelease versions having been employed in Greif19; Raaijmakers19; Raaijmakers20; Raaijmakers21; Rutherford:2022xeb.
We compare our new EOS inferences to the âNewâ scenario from rutherford_constraining_2024. This scenario is defined by the combination of the ST-U solution of PSRÂ J07406620 from salmi_radius_2024, the ST+PDT solution of PSRÂ J00300451 from vinciguerra_updated_2024, the CST+PDT solution of PSRÂ J04374715 from choudhury_nicer_2024, and the mass-tidal deformability posteriors of GW170817 and GW190425 from PhysRevX.9.011001 and abbott_gw190425_2020, respectively. We investigate the effect of adding the new mass-radius result of PSRÂ J06143329 to this data set. Interestingly, with the new result from PSRÂ J06143329 included, the EOS inference results would be very similar if only MSPs with mass and inclination priors were accounted for and PSRÂ J00300451 (ST+PDT) left out (see the Zenodo repository of zenodo for figures and samples of the inferred EOS in that case).
The EOS inference yields mass-radius posterior distributions (Figure 9) centered around 12âkm with maximum masses for both the PP and CS extensions. Our results for the radii of 1.4 and NSs, as well as the maximum mass configurations are also reported in Table 2. These inferred mass-radius posteriors are less than one standard deviation away from rutherford_constraining_2024 âNewâ data scenario, although they shift toward lower radii by m, with a minimal impact on the maximum mass for both PP and CS models. This slight shift is not only due to the addition of PSR J06143329, but also to the fact that the inferred radius of PSR J06143329 is consistent with those of PSR J00300451 (ST+PDT) and PSR J04374715.
At the 68% credible interval, the PP model posteriors show a bimodal-like structure. This is also seen in both the PP and CS models for the distribution of the radius of a NS in the insets of Figure 9. This bimodal-like structure is already present in the posterior mass-radius distributions without the inclusion of PSR J06143329, but is less pronounced for the PP model and evenly peaked below and above 12âkm for the CS model (see Figure 7 of rutherford_constraining_2024). These bimodal-like distributions are the result of larger radii (stiffer) EOS being favored by both the prior and the large measured mass of PSR J07406620 plus smaller radii (softer) EOS being favored by both PSR J04374715, PSR J06143329, GW170817 data, and, marginally, the ST+PDT contour of PSR J00300451, which overlaps with the mass-radius contours of the latter sources. Including PSR J06143329, which offers more posterior support for mass-radius samples below 12âkm, make this bimodal-like structure slightly less pronounced for the CS model and more pronounced for the PP model. This suggests that the inclusion of PSR J06143329 slightly softens the EOS. This trend is confirmed in the pressure posterior distributions at and shown in Figure 10 (see also the values in Table 2), where the pressure is pushed to slightly lower values, while still being on the stiffer end of the priors, confirming a pattern seen in Raaijmakers21 and rutherford_constraining_2024. As expected, the bimodal-like feature is also present in the pressure posterior distributions, particularly at .
Adding PSR J06143329 allows for slightly more negative difference of the radii of 2.0 and NSs, . This can be explained by a softening of the EOS, as it would reduce more prominently the posterior radius around 2.0  than at 1.4 . This effect can be seen on the lower radius branch of the bimodal-like structure for PP. Hence, the EOS slightly softening with the inclusion of PSR J06143329 is further confirmed. These values, along with other key quantities, are provided on Table 2.
| PP | CS | |
| â [km] | ||
| â [km] | ||
| â [km] | ||
| â [] | ||
| â [km] | ||
| â [g/cm3] | ||
| â [dyn/cm2] | ||
| â [dyn/cm2] | ||
| â [dyn/cm2] | ||
| â [dyn/cm2] |
âą
Notes - and are the radii of a ââ and ââNS, with . and are the mass and radius of a non-rotating, maximum mass NS, with , and being its central energy density, density, and pressure, respectively. We also give the pressure at 2, 3, and (see Figure 10 for associated distributions). For this table only, to compare with rutherford_constraining_2024, median values with equal-tailed credible intervals are given.
In addition, if instead of the Headline ST+PDT, the PDT-Uâ geometry had been used for PSR J06143329, we expect that the new EOS inferences would still have stronger support for softer EOS than the âNewâ scenario in rutherford_constraining_2024, while the reported mass-radius contours at and confidence intervals would change notably less from the âNewâ scenario than in the ST+PDTâ case, as mass-radius posteriors from the exploratory PDT-Uâ run mostly overlap the PSR J04374715 ones (Figure 3).
6.3 Caveats
During our analyses, we did not try the Concentric Single Temperature (CST) or Concentric Double Temperature (CDT) hot region models (see riley_nicer_2019). These are made of two concentric overlapping circular spots of different temperatures. The configuration of such models can be reproduced by the PDT model, where both spot centers align, which would result in a parameter (angular distance between ceding and superseding components centers) to be close to . For the PDT-Uâ model, we found for both hot regions, while for the Headline ST+PDT, the mode closest to such a configuration had . In the latter case, although is small it is only compatible with 0 at .
We also did not test any Protruding Single Temperature (PST) hot region (a PDT region whose superseding region does not emit thermal radiation). The PST configuration can be reproduced by the PDT model, which would appear in the posterior as an unconstrained distribution of the superseding component temperature, ranging from the lowest prior value up to , where the superseding region starts to emit significant amount of radiation. No such structure was found in the posteriors of the ST+PDTâ and PDT-Uâ models.
Although the PST, CST and CDT geometric models did not appear as specific solutions of a PDT region, it would, however, be interesting to test them in further analyses of this source.
This issue of model choice is central in PPM, as it affects inferred radii constraints (see for instance vinciguerra_updated_2024). Choosing an overly complex model can result in biased radius measurements with overestimated evidence, as it might be the case with PDT-U. Here, we chose for our headline result to use the ST+PDTâ model over the PDT-Uâ one, which has better evidence, although not significantly. Thorough model comparison was undertaken, but we could not definitively tell whether the PDT-Uâ model was preferred or not (see Section 5.3). The similarity and consistency of radii and geometry constraints from PDT-Uâ and ST+PDT, as well as ST+PDTâ being the simplest model fitting the data well, motivated the choice of the latter for our headline run. Our limited computation resources also weighed in on this choice. Further X-PSI analyses of this source could be carried out with a high resolution PDT-Uâ model.
Although many tests have been done in Section 5 to assess the efficiency of the sampling and the accuracy of the likelihood, it is impossible to assert the completeness of the sampling. There is, however, no strong evidence supporting the incomplete sampling case and the procedure seems to have effectively converged as the headline ST+PDTâ run yielded similar results as the exploratory run. In retrospect, this supports our choice of LR for exploratory runs (Section 3.1), which proved to be a sufficiently high resolution as we found no discrepancies between the exploratory and Headline ST+PDTâ runs. Using higher resolution X-PSI settings for the Headline ST+PDTâ was, however, required to look for complex geometries that could have been missed during sampling. These higher resolution X-PSI settings were picked judiciously so that computation time was not too heavily affected while improving accuracy by one order of magnitude (Appendix A). Moreover, we also showed that the LR exploratory run posterior samples have likelihoods consistent with that of a high resolution likelihood (Section 5.1) and that our inference is resilient to noise (Section 5.2).
6.4 Different modes and surface geometries
As the inferred distributions appear to have different modes whose distributions do not overlap at in some region of the parameter space, using the multi-modal version of MultiNest increases the acceptance rate of samples, which can help to reduce computation time while retaining sampling accuracy (shaw_efficient_2007; MultiNest_2009). The inferred radii differ for the two possible headline geometries (Figure 4), which individually have different radius values and lower uncertainties (see posterior distributions in Appendix B). Hence, breaking this multi-modality and converging on one single mode would reduce the total uncertainty in the radius measurement by a few hundred meters.
Physically motivated hot region priors can help favoring one geometry over the other by ruling out geometric configurations (Kalapotharakos_geometry_2009; Das_PPM_GRMHD_2025). Such priors can also help differentiate between ST+PDTâ and PDT-Uâ models.
Future improvements in the available data with more NICER exposure, more accurate priors on distance, inclination and mass from radio pulsar timing (e.g., with the Square Kilometer Array, ska_2025_densematter) can also help break this multi-modality. As the XMM exposure is only ks long888To put this in context with previous analyses, the MOS1/2 exposures were ks for PSRÂ J07406620 (salmi_radius_2022), ks long for PSRÂ J00300451 (vinciguerra_updated_2024), and ks long for PSRÂ J04374715 (choudhury_nicer_2024)., more XMM data would be valuable to further constrain the background, one of the main limitations of our study. This will in turn help to tighten the uncertainties on the inferred radius. Higher quality data obtained with next-generation X-ray observatories such as eXTP (, Li_eXTP_2025) and NewAthena (, cruise_newathena_2024) will also be of great help for this endeavor. These observatories will also decrease uncertainties of individual modes thanks to their large effective area, time resolution and constrained background.
7 Conclusion
In this work, we have presented the radius measurement of PSRÂ J06143329 with PPM using Bayesian inference on joint NICER and XMM data. Our analyses benefited from tight priors on both mass and inclination. The results were consistent across tested models, all revealing a similar geometry composed of polar and equatorial hot regions. The inferred radii for different models were all consistent at the 68% confidence level and the final measured radius is km for a mass of using an ST+PDTâ model. This radius range is consistent with GW observation of GW170817 as well as with PPM measurements of PSRÂ J04374715 and the ST+PDTâ model of PSRÂ J00300451, all of which are in the same mass range.
Various tests were carried out to assess the validity of different inferences. It was found that (1) the effect of LR likelihood on sampling for the exploratory runs was limited, (2) the inferred parameter posterior distributions were not affected by noise realization, and (3) model comparison disfavored ST-Uâ and did not significantly favor PDT-Uâ over ST+PDT. As inferred radii and geometries were consistent, we chose the simpler ST+PDTâ model, which explains well the data, to run the Headline result.
The effect of the tight inclination prior and atmosphere ionization state on the inferred radius and geometry for the ST+PDTâ model was also tested. These show little dependency of inferred parameters on these assumptions, although the inclination prior helped to better constrain the geometry and the radius. Along with the consistent results across all different geometric models, these make the constraints obtained in this study resilient to various hypotheses and systematics.
Further EOS analyses of the impact of this radius measurement result in a slight softening of the EOS compared to our previous work (rutherford_constraining_2024), as well as shifting the allowed mass-radius region toward lower radii by âm for both high-density EOS extensions, making them compatible with previous analyses within less than one standard deviation. For the maximum mass, this new radius constraint has little impact.
This radius measurement could be narrowed down in the future with better geometry priors to break degeneracies, with improved X-ray data from XMM, NewAthena, or eXTP to better constrain the background or with a tighter prior on the mass from radio timing.
Acknowledgments: The authors acknowledge fruitful discussion with Vivek Krishnan and Michael Kramer about the pulsar mass measurements from radio timing with MeerKAT. The authors thank the referee for their suggestions, which contributed to improve this article. This study has been partially supported through the grant EUR TESS N°ANR-18-EURE-0018 in the framework of the Programme des Investissements dâAvenir. This work was performed using HPC resources from CALMIP (Grant 2016-19056). Part of this work has been supported by the CEFIPRA grant IFC/F5904-B/2018 and ANR-20-CE31-0010 (MORPHER). All researchers from IRAP acknowledge the support of the CNES. Portions of this work performed at NRL were supported by NASA. D.C., B.D. and A.L.W. acknowledge support from ERC Consolidator grant No. 865768 AEONS (PI: Watts). M.H., Y.K. and A.L.W. acknowledge support from the NWO grant ENW-XL OCENW.XL21.XL21.038 Probing the phase diagram of Quantum Chromodynamics (PI: Watts). The work of M.M., A.S., and I.S. was supported by the European Research Council (ERC) under the European Unionâs Horizon 2020 research and innovation program (Grant Agreement No. 101020842). N.R. acknowledges support from NASA grant No.80NSSC22K0092. S.B. acknowledges funding from NASA grants 80NSSC20K0275 and 80NSSC22K0728. The work presented is based in part on observations with XMM-Newton, an ESA Science Mission with instruments and contributions directly funded by ESA Member states and NASA. This research has made use of data products and software provided by the High Energy Astrophysics Science Archive Research Center (HEASARC), which is a service of the Astrophysics Science Division at NASA/GSFC and the High Energy Astrophysics Division of the Smithsonian Astrophysical Observatory. This work has relied on NASAâs Astrophysics Data System (ADS) bibliographic services and the ArXiv.
Facilities: NICER, XMM-Newton
Softwares: ArviZ (kumar_arviz_2019), Cython (cython2011), fgivenx (fgivenx), GetDist (Lewis19, https://github.com/cmbant/getdist), GNU Scientific Library (GSL; Gough:2009), HEASoft (2014ascl.soft08004N), IPython (IPython2007), Jupyter (Kluyver:2016aa), Matplotlib (Hunter:2007; matplotlibv2), MPI (MPI), MPI for Python (mpi4py), MultiNest (MultiNest_2009), NEOST v2.1.0 (https://github.com/xpsi-group/neost; Raaijmakers24), nestcheck (higson2018nestcheck; higson2018sampling; higson2019diagnostic), NumPy (Numpy2011), OpenMP (openmp), PyMultiNest (PyMultiNest), Python/C language (python2007), SAS (2004ASPC..314..759G) SciPy (Scipy), X-PSI v2.2.7 (https://github.com/xpsi-group/xpsi; xpsi)
Appendix A Test of X-PSI resolution
To limit the computation time needed for our Headline result, the effects of the different X-PSI settings on likelihood computation time and accuracy were tested on a set of samples from the posterior of the exploratory ST+PDTâ model (see Section 4.2). Our goal is to attain higher accuracy while limiting computation time.
The different settings whose effects on likelihood accuracy were studied are given in Table 3. sqrt_num_cells is the square root of the approximate number of cells used to represent a spot and num_rays is the number of emission angle grid points for which light bending angles are precomputed. When the hot region has two temperature components (ceding and superseding), min_sqrt_num_cells and max_sqrt_num_cells are, respectively, the minimum and maximum values of sqrt_num_cells for the individual components. num_leaves and num_energies are, respectively, the sizes of the regularly spaced arrays of phase and energy over which the model is computed. Folding with the instrument response and integrating over the data phase bins, we get a modeled observation for a set of parameters, which is compared to the data to compute the likelihood. Increasing any of these values improves the accuracy of the likelihood computation.
We first investigated the effect of individual X-PSI settings, all other settings being fixed to their LR value. For varying setting values, we measured and registered the computation time along with the likelihood value for each sample in our set. These are then compared to the highest resolution setting tested. Results are plotted on the first two rows of Figure 11.
From these investigations, we can see that the computation time has a linear dependence on num_leaves and a superlinear dependence sqrt_num_cells, while it is sublinear for the other two settings. Moreover, increasing sqrt_num_cells does not improved the likelihood accuracy as much as it does for the other settings. It is hence important to keep this setting at values tailored to the problem in order to reduce computation time. For instance, ST regions need less sqrt_num_cells than PDT ones for an accurate likelihood computation, as the latter is more geometrically complex, especially in the case of thin and elongated structures. Having min_sqrt_num_cells and max_sqrt_num_cells different than sqrt_num_cells also allows for more accurate model computation with PDT geometries by using an adaptive mesh to represent the hot regions.
Different X-PSI resolutions, with different settings given in Table 3, were then investigated in a similar manner. Note that the LR settings were used for the exploratory runs and that the XR settings are different for ST and PDT. Moreover, the additional Flex settings configurations allow min_sqrt_num_cells and max_sqrt_num_cells to be different to sqrt_num_cells. Results for the different configurations can be found on the last row of Figure 11.
| Settings | LR | MR | HR | LRÂ Flex | MRÂ Flex | XR - ST | XR - PDT |
| num_rays | 100 | 512 | 640 | 100 | 512 | 512 | 512 |
| num_leaves | 32 | 64 | 128 | 32 | 64 | 64 | 64 |
| num_energies | 64 | 128 | 256 | 64 | 128 | 64 | 64 |
| sqrt_num_cells | 16 | 32 | 64 | 16 | 32 | 16 | 32 |
| min_sqrt_num_cells | 16 | 32 | 64 | 10 | 16 | 16 | 16 |
| max_sqrt_num_cells | 16 | 32 | 64 | 32 | 80 | 16 | 80 |
The LRÂ Flex configuration does not give strong accuracy improvements and is slower than the previously used LR. All Mixed, MR, and MRÂ Flex give comparable likelihood accuracies. Among these, the Mixed configuration gives computation time times faster than the others. Hence, this was used for the Headline result.
Appendix B Headline result
The Headline ST+PDTâ run summary is given in Table 4. The posterior distribution of a subset of parameters from the headline ST+PDTârun, with individual modes detailed, is given in Figure 12. The figure of the full posterior distributions, including all parameters, can be found in the Zenodo repository (zenodo).
| Parameter | Description | Prior PDF | |||
| [ms] | Coordinate spin period | fixed | |||
| [] | Gravitational mass | 1.548 | 0.016 | ||
| [km] | Coordinate equatorial radius | 12.61 | 1.330 | ||
| [kpc] | Earth distance | 0.6685 | 0.064 | ||
| Sine of angle from line of sight to spin axis | 0.999575 | 0.004 | |||
| Compactness condition | |||||
| Surface gravity condition | |||||
| [cycles] | p region - initial phase | 0.412 | 5.213 | ||
| [rad] | p region - center colatitude | 1.424 | 0.887 | ||
| [rad] | p region - angular radius 1 | 0.067 | 3.271 | ||
| p region - effective temperature | 6.044 | 3.972 | |||
| [cycles] | s region - initial phase a | 0.919 | 2.119 | ||
| [rad] | s region, superseding - center colatitude | 0.191 | 2.254 | ||
| [rad] | s region, superseding - angular radius | 0.051 | 1.115 | ||
| s region, superseding - effective temperature | 6.209 | 1.911 | |||
| [rad] | s region, ceding - center colatitude | 0.10 | 2.665 | ||
| [rad] | s region, ceding - angular radius | 0.56 | 0.385 | ||
| [rad] | Azimuthal offset between the s components | 2.05 | 2.519 | ||
| s region, ceding - effective temperature | 5.73 | 1.846 | |||
| Non overlapping hot regions condition | |||||
| Interstellar neutral Hydrogen column density | 0.124 | 3.064 | |||
| NICER effective area scaling factor | 1.17 | 0.071 | |||
| PN effective area scaling factor | 0.867 | 0.3230 | |||
| MOS1 effective area scaling factor | 0.982 | 0.292 | |||
| MOS2 effective area scaling factor | 0.911 | 0.142 | |||
| Sampling information | |||||
| Number of parameters: | |||||
| Number of live points: | |||||
| Sampling efficiency: | |||||
| Evidence tolerance: | |||||
| Likelihood evaluations: | |||||
| Estimated evidence: | |||||
| Computation time: CPU hours | |||||
| on Intel Skylake 6140 at 2.3 Ghz with 36 cores |
âą
Notes - We report the prior PDFs, the median with the surrounding credible intervals , the maximum likelihood
parameter vector , and the KullbackâLeibler divergence in bits representing prior-to-posterior information gain. -
âą
a Unlike most previous analyses, the secondary spot phase is not shifted by cycles.